OpenStudy (anonymous):

is f(x)=4^3+3x odd or niether

OpenStudy (anonymous):

hint what are the exponents?

OpenStudy (anonymous):

btw i am assuming this is actually \[f(x)=4x^3+3x\] am i right?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

how is f(-x) related to f(x)? if f(-x)=f(x) the function is even. if f(-x)=-f(x) the function is odd. otherwise the function is neither even nor odd

OpenStudy (anonymous):

so is the answer neither?

OpenStudy (unklerhaukus):

what do you get for f(-x) ?

OpenStudy (anonymous):

what are the exponents?

OpenStudy (anonymous):

i.e. what are the degrees of each term?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

\[\large f(x)=4x^{\color{red}3}+x^{\color{blue}1}\]

OpenStudy (anonymous):

is it odd or neither

OpenStudy (unklerhaukus):

f(-x) = 4(-x)^3+(-x) =